OpenStudy (fibonaccichick666):

I just need someone to tell me what this question wants in dummy terms: Suppose that p is a polynomial of degree d. Given any \(x_0\in R\), show that the \((d+1){th}\) Taylor Polynomial for p at \(x_0\) is equal to p.

OpenStudy (fibonaccichick666):

I don't understand how you can have a (d+1)th for a d degree function.

OpenStudy (fibonaccichick666):

@SithsAndGiggles

Miracrown (miracrown):

d+1 degree Taylor polynomial, do you know what a taylor polynomial is?

ganeshie8 (ganeshie8):

i think you need to show that after differentiating a polynomial d+1 times, you get a zero term and all remaining terms will be 0

OpenStudy (fibonaccichick666):

yes, it's more of a the d+1 will end with a 0. because past d the deriv is zero. So I don't understand what it's asking

OpenStudy (fibonaccichick666):

oh, thanks ganeshie. I guess my reason for confusion is what I am supposed to show

Miracrown (miracrown):

what it is saying if we plug in x0 for x in the polynomial of degree d it should equal the same thing as a taylor series of degree d+1

ganeshie8 (ganeshie8):

yeah i have no clue yet how to start the proof too, but it is clear what we need to prove

OpenStudy (fibonaccichick666):

I know how to prove it, but I just didn't understand the problem thanks

OpenStudy (fibonaccichick666):

can you medal Miracrown for helping for me ganeshie?

OpenStudy (dan815):

no medal me

ganeshie8 (ganeshie8):

already gave it to mira :) fib you medal dan maybe..

OpenStudy (dan815):

<3

Miracrown (miracrown):

and x0 is a real number

OpenStudy (dannyrod2000):

i medal ganeshi

OpenStudy (dan815):

ah basically its a matter of taylor series expansion being the approimations wrt to derivatives

OpenStudy (dan815):

for a certain nth degree polynomial all the succeding taylor approximations after d+1 yeild zero so

ganeshie8 (ganeshie8):

how to start a proof like this

ganeshie8 (ganeshie8):

it should be sufficient to just state it in words right ?

OpenStudy (fibonaccichick666):

i am going with based off of the definition of the taylor p, if a deriv =0 then the term of the poly=0 so if we show d+1,2,3 by induction =0 then we are left with the taylor of d

OpenStudy (dan815):

all polynomials can be rewritten in form kth polynomial = ax^k+bx^(k-1).... Therefore the k+1 derivative of this kth polynomial = 0

OpenStudy (dan815):

(k+1)th derivative

ganeshie8 (ganeshie8):

consider a polynomial of degree "d" by taylor theorem, the taylor series equals this polynomial. however after differentating this polynomial for d+1 times, we get f^(d+1) = 0 and thus all the subsequent will be 0. that means the first d+1 terms of taylor series equals the polynomial \(\square\)

OpenStudy (fibonaccichick666):

y'all want some more? I have like 4 more due tomorrow at 1 that I need interpretations so I understand what it is asking

OpenStudy (dan815):

ya only if u gimme ur medals this time

OpenStudy (fibonaccichick666):

but you don't have to complete them like you are

OpenStudy (dan815):

bahahahah

OpenStudy (fibonaccichick666):

as told by ganeshie, I have passed the medal

OpenStudy (fibonaccichick666):

Also, I hate analysis. Any help is appreciated

ganeshie8 (ganeshie8):

just a quick q, does taylor series exists for all functions ?

OpenStudy (fibonaccichick666):

well, essentially it is the MVT for derivatives, so so long as it's differentiable, I think so

ganeshie8 (ganeshie8):

consider a polynomial of degree "d" by taylor theorem, the taylor series, `if it exists`, equals this polynomial. however after differentating this polynomial for d+1 times, we get f^(d+1) = 0 and thus all the subsequent will be 0. that means the first d+1 terms of taylor series equals the polynomial

ganeshie8 (ganeshie8):

yeah its redundant if we agree that the series exists for all polynomials

ganeshie8 (ganeshie8):

as all polynomials are differentiable

OpenStudy (fibonaccichick666):

def exists for all poly, but what about trancendentals? (sp? sorry)

OpenStudy (dan815):

what does a transcendental mean

OpenStudy (fibonaccichick666):

A transcendental function is a function that does not satisfy a polynomial equation whose coefficients are themselves roots of polynomials, in contrast to an algebraic function, which does satisfy such an equation.

OpenStudy (fibonaccichick666):

hey it's essentially p=np.... freaky

OpenStudy (dan815):

dp/dn= p+dp/dn D_2 = dp/dp + d^2p/dn^2 D_N= d^(n-1)/dp^(n-1) +d^np/dn^n

OpenStudy (dan815):

all derivatives exist

OpenStudy (dan815):

so ud need infinite taylor approximations

OpenStudy (fibonaccichick666):

yep like sin and cos

OpenStudy (dan815):

yah

OpenStudy (dan815):

sin and cos are infinite polynomials kinda

OpenStudy (dan815):

like u know ax^4+bx^3... looks like this w